OpenStudy (mtalhahassan2):

An object falls from rest 100m above the ground. If the acceleration due to gravity is 10m/s/s[down], a) determine the time it takes to hit the ground, b) the velocity at the instant it impacts the ground, c) prove the equations...

OpenStudy (shamim):

a)

OpenStudy (shamim):

Required formula is h=ut+.5gt^2

OpenStudy (shamim):

b) v^2=u^2+2*gh

OpenStudy (mtalhahassan2):

hy

OpenStudy (shamim):

Plz feel free to ask for more clarification

OpenStudy (shamim):

Given u=0 g=10m/s h=100m

OpenStudy (shamim):

a) t=?

OpenStudy (shamim):

b)v=?

OpenStudy (anonymous):

i believe your supposed to clarify and explain how to get the answer, not just give the answer